2x+5y=14
2x-2y=-36

Found 2 solutions by lynnlo, Alan3354:
Answer by lynnlo(4176)

(1)x=7-5y/2
(2)x=y-18

Answer by Alan3354(69443)

2x+5y=14
2x-2y=-36
-------------- Subtract
7y = 50
y = 50/7
----------
2x-2y=-36
x - y = -18
x - 50/7 = -18
x = 76/7